An interest rate of 14,90% per year,

compounded every 3 months. What is equivalent interest rate compounded weekly?


1
Expert's answer
2022-03-15T17:22:26-0400

"A=P(1+\frac{r}{n})^{nt}"

Let P = 1, t = 1

"A=1(1+\frac{0.1940}{4})^4\\A=1.1575340471"

Compounded every week

"365÷7\approx52 \space weeks\\n=52\\1.1575340471=(1+\frac{r}{52})^{52}"

"\sqrt[52]{1.1575349471}=(1+\frac{r}{52})\\52\times \sqrt[52]{1.1575349471}-52=r\\r=0.14649789"


equivalent interest rate compounded weekly

r=0.14649789